# Physics-Informed Transformer operator for the prediction of three-dimensional turbulence

### v3

@@ III Numerical results
III.1 Decaying homogeneous isotropic turbulence at statistically stationary initial condition
The direct numerical simulation of decaying homogeneous isotropic turbulence (HIT) is conducted in a $(2\pi)^{3}$ cubic domain under periodic boundary conditions, employing a uniform grid resolution of $128^{3}$.[96] The pseudo-spectral method is applied for spatial discretization of the governing equations, coupled with a second-order Adams-Bashforth scheme for time integration.[5, 59] In addition, high-wavenumber Fourier modes are truncated in accordance with the two-thirds rule to suppress aliasing.
The statistically stationary initial condition for the decay HIT is obtained from a fully developed turbulent flow sustained by external forcing. In order to maintain a stationary turbulent velocity field, large-scale forcing is applied by holding the velocity spectrum in the two lowest wavenumber shells fixed at the values $E_{0}(1)=1.242477$ and $E_{0}(2)=0.391356$. The kinematic viscosity is set to $\nu=0.015625$ and the corresponding Taylor Reynolds number is $Re_{\lambda}\approx 60$. The DNS time step $\Delta t$ is set to 0.001. The simulation is initialized with a random velocity field following a Gaussian distribution in spectral space. We maintain 10,000 time steps using the forcing term until the flow becomes statistically steady. Thereafter, the forcing term is removed, allowing the turbulence to decay over 5000 time steps, enabling the development of isotropic decay turbulence.[96]
The DNS data are filtered to obtain large-scale flow fields with the sharp spectral filter. The filtering is performed on a $32^{3}$ grid with a cutoff wavenumber of $k_{c}=10$.[96] Here, we generated 5,000 stationary initial turbulent velocity fields for training and 20 samples for testing. We save the numerical solution for each time node at every 20 DNS time steps. To evaluate the performance and generalization ability of the model, the a posteriori analysis employs five independent benchmark cases with different initial conditions, where each simulates a complete 5000-step decay period of homogeneous isotropic turbulence. We use the high-fidelity fDNS data for both training and testing. The data are formatted as tensors of size $N_{t}$,32,32,32,3], where $N_{t}$ is the number of input time steps. The physics-informed models are trained on a dataset of 5000 samples, each containing one ($N_{t}=1$) velocity field at the time $t=10000\Delta t$. The test set consists of 20 additional samples. Each sample is a sequence of length $N_{t}=11$, containing velocity fields from $t=10,000\Delta t$ to $t=10,200\Delta t$. The detailed simulation parameters are provided in Table 1.[96]
@@ III.1 Decaying homogeneous isotropic turbulence at statistically stationary initial condition
The training hyperparameters for all data-driven models are summarized in Table 2. The decay period values (4000, 10000, 20000) represent a learning rate schedule.[96] The learning rate is multiplied by the decay factor at these designated epochs to refine the optimization process. Here, we utilize the Adam optimizer for training and the GELU function for activation.[31, 23] Physics-Informed models minimize the PDE loss defined by (3) and (4).
@@
Machine learning approaches for numerical simulations are distinguished by their lower computational costs. Table 4 compares the computational efficiency of the traditional model SM with three physics-informed models. Here, the physics-informed models are trained and tested on an NVIDIA A100 PCIe 40GB graphics processing unit (GPU) paired with an Intel 6248R CPU @3.0GHz, while the SM model is simulated on an Intel Xeon Gold 6148 CPU (16 cores @ 2.40 GHz).
@@
III.2 Decaying homogeneous isotropic turbulence at random initial condition
@@
III.3 Forced homogeneous isotropic turbulence
@@
IV Effect of Hyperparameters on Model Performance
@@
V Automatic Optimization Of Subgrid-Scale Coefficient
In the Smagorinsky model for large eddy simulation (LES), the coefficient $C_{\text{smag}}$ is typically treated as an empirical constant. However, the optimal value of this coefficient is often unknown for different flow conditions. Leveraging the physics-informed architecture, we can embed $C_{\text{smag}}$ as a trainable parameter within the neural network, allowing the model to automatically identify the optimal coefficient value and minimize the PDE loss.[96]
We generate datasets in decaying HIT at stationary initial condition for learning the $C_{\text{smag}}$ coefficient, which consists of a single set of either SM or fDNS dataset. These datasets are respectively derived from LES simulations employing the Smagorinsky model ($C_{\text{smag}}^{2}=0.01$) and DNS. It should be emphasized that these datasets are used only for updating the $C_{\text{smag}}$ coefficient.[96] The SM and fDNS datasets are generated over the temporal interval from $t=10,000\Delta t$ to $10,200\Delta t$ collected every $20\Delta t$.
During training, the model takes original fDNS dataset and additional SM or fDNS datasets as inputs. Two independent Adam optimizers are employed: one for the neural operator and another for the Smagorinsky coefficient $C_{\text{smag}}$. The output velocity fields and $C_{\text{smag}}$ are then both used to compute $\mathcal{L}_{PDE}$. In this method, the Loss is constructed as follows:[96]
@@ V Automatic Optimization Of Subgrid-Scale Coefficient
With the parameter set to $\gamma=50$,[96] we evaluate the automatic learning procedure of the Smagorinsky coefficient using PITO model in the decaying HIT at stationary initial condition. The evolutions of the loss curve and the automatic learning of the coefficient $C_{\text{smag}}$ are shown in Figs. 20(a) and (b). The Smagorinsky coefficient $C_{\text{smag}}$ reaches final converged values of 0.0969 and 0.1871 for the SM and fDNS datasets, respectively. Furthermore, in the a posteriori tests in Figs. 20(c) and (d), the model demonstrates an excellent agreement with the fDNS results, performing as effectively as the original model.
@@ Appendix A Fourier Neural Operator (FNO)
The Fourier neural operator (FNO) is a neural operator that parameterizes the integral kernel in the spectral domain.[39] The FNO architecture is shown in Fig. 22. In the FNO architecture, each Fourier layer updates the latent feature $v_{l}$ by combining a local linear transformation with a spectral convolution:

### v4

@@ 3 Numerical results
3.1 Decaying homogeneous isotropic turbulence at statistically stationary initial condition
The direct numerical simulation of decaying homogeneous isotropic turbulence (HIT) is conducted in a $(2\pi)^{3}$ cubic domain under periodic boundary conditions, employing a uniform grid resolution of $128^{3}$ [90]. The pseudo-spectral method is applied for spatial discretization of the governing equations, coupled with a second-order Adams-Bashforth scheme for time integration [96, 97]. In addition, high-wavenumber Fourier modes are truncated in accordance with the two-thirds rule to suppress aliasing.
The statistically stationary initial condition for the decay HIT is obtained from a fully developed turbulent flow sustained by external forcing. In order to maintain a stationary turbulent velocity field, large-scale forcing is applied by holding the velocity spectrum in the two lowest wavenumber shells fixed at the values $E_{0}(1)=1.242477$ and $E_{0}(2)=0.391356$. The kinematic viscosity is set to $\nu=0.015625$ and the corresponding Taylor Reynolds number is $Re_{\lambda}\approx 60$. The DNS time step $\Delta t$ is set to 0.001. The simulation is initialized with a random velocity field following a Gaussian distribution in spectral space. We maintain 10,000 time steps using the forcing term until the flow becomes statistically steady. Thereafter, the forcing term is removed, allowing the turbulence to decay over 5000 time steps, enabling the development of isotropic decay turbulence [90].
The DNS data are filtered to obtain large-scale flow fields with the sharp spectral filter. The filtering is performed on a $32^{3}$ grid with a cutoff wavenumber of $k_{c}=10$ [90]. Here, we generated 5,000 stationary initial turbulent velocity fields for training and 20 samples for testing. We save the numerical solution for each time node at every 20 DNS time steps. To evaluate the performance and generalization ability of the model, the a posteriori analysis employs five independent benchmark cases with different initial conditions, where each simulates a complete 5000-step decay period of homogeneous isotropic turbulence. We use the high-fidelity fDNS data for both training and testing. The data are formatted as tensors of size $N_{t}$,32,32,32,3], where $N_{t}$ is the number of input time steps. The physics-informed models are trained on a dataset of 5000 samples, each containing one ($N_{t}=1$) velocity field at the time $t=10000\Delta t$. The test set consists of 20 additional samples. Each sample is a sequence of length $N_{t}=11$, containing velocity fields from $t=10,000\Delta t$ to $t=10,200\Delta t$. The detailed simulation parameters are provided in Table 1 [90].
@@ 3.1 Decaying homogeneous isotropic turbulence at statistically stationary initial condition
The training hyperparameters for all data-driven models are summarized in Table 2. The decay period values (4000, 10000, 20000) represent a learning rate schedule [90]. The learning rate is multiplied by the decay factor at these designated epochs to refine the optimization process. Here, we utilize the Adam optimizer for training and the GELU function for activation [98, 99]. Physics-Informed models minimize the PDE loss defined by (3) and (4).
@@
Machine learning approaches for numerical simulations are distinguished by their lower computational costs. Table 4 compares the computational efficiency of the traditional model SM with three physics-informed models. Here, the physics-informed models are trained and tested on an NVIDIA A100 PCIe 40GB graphics processing unit (GPU) paired with an Intel 6248R central processing unit (CPU) @3.0GHz, while the SM model is simulated on an Intel Xeon Gold 6148 CPU (16 cores @ 2.40 GHz).
@@
3.2 Decaying homogeneous isotropic turbulence at random initial condition
@@
3.3 Forced homogeneous isotropic turbulence
@@
4 Effect of Hyperparameters on Model Performance
@@
5 Automatic Optimization Of Subgrid-Scale Coefficient
In the Smagorinsky model for large eddy simulation (LES), the coefficient $C_{\text{smag}}$ is typically treated as an empirical constant. However, the optimal value of this coefficient is often unknown for different flow conditions. Leveraging the physics-informed architecture, we can embed $C_{\text{smag}}$ as a trainable parameter within the neural network, allowing the model to automatically identify the optimal coefficient value and minimize the PDE loss [90].
We generate datasets in decaying HIT at stationary initial condition for learning the $C_{\text{smag}}$ coefficient, which consists of a single set of either SM or fDNS dataset. These datasets are respectively derived from LES simulations employing the Smagorinsky model ($C_{\text{smag}}^{2}=0.01$) and DNS. It should be emphasized that these datasets are used only for updating the $C_{\text{smag}}$ coefficient [90]. The SM and fDNS datasets are generated over the temporal interval from $t=10,000\Delta t$ to $10,200\Delta t$ collected every $20\Delta t$.
During training, the model takes original fDNS dataset and additional SM or fDNS datasets as inputs. Two independent Adam optimizers are employed: one for the neural operator and another for the Smagorinsky coefficient $C_{\text{smag}}$. The output velocity fields and $C_{\text{smag}}$ are then both used to compute $\mathcal{L}_{PDE}$. In this method, the Loss is constructed as follows [90]:
@@ 5 Automatic Optimization Of Subgrid-Scale Coefficient
With the parameter set to $\gamma=50$ [90], we evaluate the automatic learning procedure of the Smagorinsky coefficient using PITO model in the decaying HIT at stationary initial condition. The evolutions of the loss curve and the automatic learning of the coefficient $C_{\text{smag}}$ are shown in Figs. 20(a) and (b). The Smagorinsky coefficient $C_{\text{smag}}$ reaches final converged values of 0.0969 and 0.1871 for the SM and fDNS datasets, respectively. Furthermore, in the a posteriori tests in Figs. 20(c) and (d), the model demonstrates an excellent agreement with the fDNS results, performing as effectively as the original model.
@@ Appendix A Fourier Neural Operator (FNO)
The Fourier neural operator (FNO) is a neural operator that parameterizes the integral kernel in the spectral domain [42]. The FNO architecture is shown in Fig. 22. In the FNO architecture, each Fourier layer updates the latent feature $v_{l}$ by combining a local linear transformation with a spectral convolution: